OpenStudy (anonymous):

Write the equation for the give points on a graph y8, x -5, y-8

OpenStudy (anonymous):

does y8 mean that the y intercept is 8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In that case, use the for y = mx+b (where m is the slope and b is the y int) so, we have: y = mx + 8 to get m, sub in the point (-5,-8) -8 =m(-5) +8 -16 = -5m m = 16/5 which now gives us: y = (16/5)x +8

OpenStudy (anonymous):

ok thank u